Two trains start from the same station and run at an opposite directions at an average rate of 64 kilometers per hour and 86 kilometers per hour, respectively. In how many hours will they be 225 kilometer apart?
Answer by mananth(16946)

two trains
one train 64 kmph
other train 86 kmph
..
they are going away from each other and co-linear.
so the effective speed = 64+86 = 150 kmph
distance = 225 km
...
time = distance/speed = 225/150
time = 1.5 hours they will be 225 km apart
...
m.ananth@hotmail.ca